# Membership problems in braid groups and Artin groups

### v2

@@ 1. Introduction
Algorithmic problems are of central importance in the study of braid groups, and more generally of Artin groups. For braid groups, the word and conjugacy problems are both decidable, and they have been studied extensively from many different viewpoints in numerous important papers; see e.g. [Art47, BKL98, Deh97, EM94, Gar69, GKT02, GMV14]. For more background on braid groups, including the history and motivation for their study, including connections with representation theory, algebraic geometry and topology, and mathematical physics, we refer the reader to the book [KT08].
Artin groups, which are a natural generalization of braid groups closely related to Coxeter groups, first appeared in the literature in the papers [Del72, BS72]. Even after fifty years of intensive study, Artin groups remain a highly mysterious class and in general very little is known about the algorithmic properties of Artin groups; see e.g. the survey articles [Bir74, GP12] and [McC17]. In particular, both the word and conjugacy problems remain open in general for Artin groups. A lot of work has been done in this area, and it continues to be a highly active area of research. Finite type Artin groups (which includes all braid groups) have nice algorithmic properties e.g. they are biautomatic [Cha92, Cha95] which gives efficient solutions to the word and conjugacy problems in these cases. Other papers on Artin groups and their algorithmic properties include e.g. [BM00, CGW09, HH23, HM99, HO20, HR15, HR13, HR12, MS17].
@@
In this article we will study the membership problem in finitely generated submonoids, and more generally rational subsets, of Artin groups. We will also investigate a range of other related membership-type decision problems including the fixed-target submonoid membership problem and semigroup intersection problem. Further motivation for studying the submonoid membership problem in Artin groups comes from the fact that the word problem for Artin groups can be reduced to deciding membership in particular finitely generated submonoids. For example, it is not hard to see that for any Artin group if membership in the Artin submonoid of that Artin group is decidable then it follows that the Artin group has decidable word problem. Indeed, in this case to decide whether a given word equals the identity one can first check whether it belongs to the Artin submonoid and, if it does, use the fact that the Artin monoid (trivially) has decidable word problem to test whether the word equals the identity in the Artin monoid, and hence in the group.
@@
In addition to the submonoid membership problem, there are a range of other natural membership-type problems in groups that have been introduced and investigated including: rational subset membership, semigroup intersection, the group problem, and the identity problem (see below for definitions of these notions). The study of these decision problems has motivation coming from both mathematics and theoretical computer science, and is a burgeoning field of research with a lot of recent intensive activity; see e.g. the recent survey articles [Don23, Loh24] and also [Loh15]. For a helpful diagram showing how all these decision problems relate to each other we refer the reader to [Don23, Fig. 1]. In addition to the submonoid membership problem, we shall also consider these other decision problems listed above for Artin groups. We will show that the our classification of Artin groups with decidable submonoid membership problem also gives a classification of Artin groups with decidable rational subset membership problem, and with decidable semigroup intersection problem; see Corollary 4.6.
@@
We do not currently know if this extends to also classify the Artin groups with decidable group problem or identity problem. In this direction in Theorem 4.5 we identify some families of Artin groups for which the identity problem and group problem are undecidable, and in the remaining cases of our classification we show that a related problem is undecidable: the fixed-target submonoid membership problem (that we shall define). These results are established by proving several new undecidability results for the group $A(P_{4})$ in Section 3.
@@
In addition to presenting these algorithmic results on braid and Artin groups, another aim of this paper is to bring to the attention to those working on group and semigroup membership problems this method of embedding the $A(P_{4})$ as an approach to proving undecidability results for groups that are known not to embed $F_{2}\times F_{2}$. Indeed, it was recently proved in [FGNB24, Corollary 6.4] that just embedding the trace monoid $T(P_{4})$ into a group is enough to yield undecidability of membership in rational subsets. We believe that there are likely many more natural situations when pairs of words cannot be embedded into a group, but $A(P_{4})$ or $T(P_{4})$ can, giving rise to undecidability results. This is, for example, exactly what happens in the case of one-relator groups and monoids; see [Gra20, FGNB24] and also the related work [MV24].
This article is structured as follows. In Section 2 we give the necessary background definitions and results on finitely generated groups and monoids, decision problems, and Artin groups including braid groups and right-angled Artin groups. In Section 3 we complete the classification of braid groups with decidable submonoid (and rational subset) membership problem initiated by Potapov in [Pot13]; see Theorem 3.1. We go on to prove several undecidability results for the right-angled Artin group $A(P_{4})$ and apply these to establish further undecidability groups for the four-strand braid group (see Theorem 3.6). Then in Section 4 we turn our attention to the more general class of Artin groups. Specifically we classify (in Theorem 4.4) the Artin groups with decidable submonoid membership problem in terms of the non-existence of certain forbidden induced subgraphs of the defining graph. We will show that the our classification of Artin groups with decidable submonoid membership problem also gives a classification of Artin groups with decidable rational subset membership problem, and with decidable semigroup intersection problem; see Corollary 4.6. In Theorem 4.5 we identify some families of Artin groups for which the identity problem and group problem are undecidable, and in the remaining cases of our classification we show that the related fixed-target submonoid membership problem is undecidable.
@@ 2.1. Monoid presentations and rational subsets
[…]
@@ 2.2. Membership problems
There are many actively studied decision problems which fall under these umbrella terms. Let $G$ be a group generated by a finite set $A$, with a surjective homomorphism $\pi\colon A^{\ast}\to G$. The (uniform) submonoid membership problem for $G$ asks, on input a finite set $X\subseteq A^{\ast}$ of words, and a word $w\in A^{\ast}$, whether or not $\pi(w)\in\operatorname{Mon}(X)$. The non-uniform submonoid membership is the same problem, except the finite subset $X$ is considered fixed and not part of the input; similarly, the fixed-target submonoid membership problem fixes the word $w$, and only has $X$ as part of the input. This problem will thus be referred to as the “fixed-target submonoid membership problem for $w$ in $M$”. All the above terminology is retained if instead of submonoids we consider membership in subsemigroups $\operatorname{Sgp}(X)$. A particular case of the fixed-target subsemigroup problem is when the fixed target is $1$, i.e. when we are given a set $X$ and asked to decide if $1\in\operatorname{Sgp}(X)$. This problem is known in the literature, e.g. [Don23], as the identity problem [^1] The reader is warned that the terminology “identity problem” is somewhat unfortunate, as this is also the classical name for the word problem, used e.g. by Dehn [Deh11].. Analogously, the (uniform) rational subset membership problem for $G$ asks, on input a finite automaton $\operatorname{\mathcal{A}}$ over $A$ and a word $w\in A^{\ast}$, whether or not $\pi(w)\in\pi(L(\operatorname{\mathcal{A}}))$. The non-uniform version of this problem is the same, except the automaton $\operatorname{\mathcal{A}}$ is fixed, and not part of the input. In principle, one could consider the fixed-target rational subset membership problem, fixing a target element $w$, and leaving $\operatorname{\mathcal{A}}$ as the only input. However, this can easily seen to be equivalent to the (uniform) rational subset membership problem; we thank C. Bodart for pointing out this fact to us.
@@
Another important membership-type problem that will be studied herein is the subsemigroup intersection problem, which asks, on input two finite subsets $X_{1},X_{2}\subseteq A^{\ast}$, whether or not $\operatorname{Sgp}(X_{1})\cap\operatorname{Sgp}(X_{2})=\varnothing$ in $G$. This also has more and less uniform, as well as fixed-target, variants. For example, one might fix one of the subsets $X_{i}$ (making the problem less uniform), or ask if some fixed element sits in the intersection (making the problem of a fixed-target type). We remark that if a group $G$ has decidable rational subset membership problem, then we can also solve the identity problem and the subsemigroup intersection problem in $G$. Indeed, the first is immediate, and the latter follows from the fact that $S_{1}S_{2}^{-1}$ is a rational subset of $G$ for any two finitely generated subsemigroups $S_{1},S_{2}$. But $S_{1}\cap S_{2}=\varnothing$ if and only if $1\in S_{1}S_{2}^{-1}$, and since we decide the latter, we can also decide the former. Finally, another decision problem is the group problem, which takes as input a finite set $X$ of words and decides whether or not $\operatorname{Sgp}(X)$ is a subgroup of $G$.
@@
The decision problems discussed in this section behave well when passing from a finitely generated group to a finitely generated subgroup. More precisely, suppose that $\mathcal{P}$ is any of the following decision problems: submonoid membership problem, subgroup membership problem, rational subset membership problem, fixed-target submonoid membership problem with a fixed target element $\gamma$, subsemigroup intersection problem, group problem, and identity problem. Then it is straightforward to see if $\mathcal{P}$ is decidable in $G$ then $\mathcal{P}$ is also decidable in $H$. This is an important tool for showing various problems are undecidable in a group, by embedding a suitable group for which we know the problem is undecidable.
@@ 2.3. Artin and braid groups
$\underbrace{abab\ldots}_{\mbox{$m$-factors}}=\underbrace{ababa\ldots}_{\mbox{$m$-factors}}$
@@
with the first set of relations being taken for all $11\leq i<n-$ 1, and the second set only being taken for those $i,j$ such that $|i-j|\geq 2$. There is a natural pictorial description of any element (braid) of a braid group, which also gives rise to a simple way to multiply braids together; see [KT08, §1.2.2]. Clearly the $n$-strand braid groups $\mathbf{B}_{n}$ are all Artin groups. For example, the braid group $\mathbf{B}_{4}$ is the Artin group $A(\Gamma)$ where $\Gamma$ is a triangle with edges labelled $(2,3,3)$.
@@ 3. Braid groups and right-angled Artin groups
Theorem 3.1 will be proved by applying several results from the literature on right-angled Artin groups and their embeddings into braid groups. Firstly we have the following result of Lohrey & Steinberg in which they classify the right-angled Artin groups with decidable submonoid, and rational subset, membership problem.
@@
A right-angled Artin group $A(\Gamma)$ has decidable submonoid membership problem if and only $\Gamma$ does not embed $C_{4}$ or $P_{4}$ as an induced subgraph. Furthermore, if $\Gamma$ does not embed $C_{4}$ or $P_{4}$, then $A(\Gamma)$ has decidable rational subset membership problem, while if $\Gamma$ does embed $C_{4}$ or $P_{4}$ then $A(\Gamma)$ contains a fixed finitely generated submonoid in which membership is undecidable.
As explained in the introduction above the group $A(C_{4})\cong F_{2}\times F_{2}$ does not embed in the braid group $\mathbf{B}_{4}$. However, the following result of Droms, Lewin & Servatius shows that the group $A(P_{4})$ does embed in $\mathbf{B}_{4}$.
@@ 3.1. A connection between rational subsets and submonoids
be a set in bijective correspondence with $Q$ where the bijection is fixed such that $q_{0}\mapsto x$ and $q_{f}\mapsto y$. For $q\in Q$, let $\widetilde{q}$ denote the image of $q$ in $\widetilde{Q}$. Clearly, $\widetilde{Q}$ generates a free subgroup of $F$ of rank $|Q|$.
@@
[…]
@@
[…]
@@
$\pi(xy^{-1})$ $=\pi(xwy^{-1})=\pi\left(\widetilde{q_{0}}(\sigma_{1}\sigma_{2}\cdots\sigma_{n})\widetilde{q_{f}}\right)$
@@
Let us now prove the equivalence (4), which will complete the proof of part (ii) of the lemma, and thereby also the entire lemma. If $\pi(xy^{-1})\in\operatorname{Mon}(\Delta)$, then clearly as $xy^{-1}\neq 1$ in $G\ast F$, we have that $\pi(xy^{-1}\in\operatorname{Sgp}(\Delta)$, and hence $\operatorname{Sgp}(xy^{-1})\cap\operatorname{Sgp}(\Delta)\neq\varnothing$ as required. Suppose next, for the converse, that $\operatorname{Sgp}(\Delta)\cap\operatorname{Sgp}(xy^{-1})\neq\varnothing$. Then some power of $xy^{-1}$ lies in $\operatorname{Sgp}(\Delta)$, i.e.
@@
in $G\ast F$, where the $p_{i}\xrightarrow{a_{i}}q_{i}$ are paths in $\operatorname{\mathcal{A}}$. By the above claim, we have $1\in h(L(\mathcal{A}))$. Hence, by (3), we have that $\pi(xy^{-1})\in\operatorname{Mon}(\Delta)$. Since $xy^{-1}\neq 1$, we have that $xy^{-1}$ lies in $\operatorname{Sgp}(\Delta)$, as required. This completes the proof of (4), and hence also the proof of the lemma. ∎

### v3

@@ 1. Introduction
Algorithmic problems are of central importance in the study of braid groups, and more generally of Artin groups. For braid groups, the word and conjugacy problems are both decidable, and they have been studied extensively from many different viewpoints in numerous important papers; see e.g. [Art47, BKL98a, Deh97, EM94, Gar69, GKT02, GMV14]. For more background on braid groups, including the history and motivation for their study, including connections with representation theory, algebraic geometry and topology, and mathematical physics, we refer the reader to the book [KT08].
Artin groups, which are a natural generalization of braid groups closely related to Coxeter groups, first appeared in the literature in the papers [Del72, BS72], in which the word problem is also considered. Even after fifty years of intensive study, Artin groups remain a highly mysterious class and in general very little is known about the algorithmic properties of Artin groups; see e.g. the survey articles [Bir74, GP12] and [McC17]. In particular, both the word and conjugacy problems remain open in general for Artin groups. A lot of work has been done in this area, and it continues to be a highly active area of research. It is also known that finite type Artin groups (which includes all braid groups) have nice algorithmic properties, and in particular have decidable word and conjugacy problem [Del72, BS72]; it is also known that they are biautomatic [Cha92, Cha95] which gives very efficient solutions to the word and conjugacy problems in these cases. Additional work on the word and conjugacy problems in finite type Artin groups include [BKL98b, BGGM07a, BGGM08, BGGM07b, GGM10b, GGM10a]. Other papers on Artin groups and their algorithmic properties include e.g. [BM00, CGW09, HH23, HM99, HO20, HR15, HR13, HR12, MS17, Pic01]; for some recent progress, we also refer the reader to [BGP22, BGCH+24, BGCMW22] and references therein.
@@
In this article we will study the membership problem in finitely generated submonoids, and more generally rational subsets, of Artin groups. We will also investigate a range of other related membership-type decision problems including the fixed-target submonoid membership problem and semigroup intersection problem. Further motivation for studying the submonoid membership problem in Artin groups comes from the fact that the word problem for Artin groups can be reduced to deciding membership in particular finitely generated submonoids. For example, it is not hard to see that for any Artin group if membership in the Artin submonoid of that Artin group is decidable then it follows that the Artin group has decidable word problem. Indeed, in this case to decide whether a given word equals the identity one can first check whether it belongs to the Artin submonoid. If it does not belong to the submonoid, then it is not equal to the identity element; and if it does, then we can use the fact that the Artin monoid (trivially) has decidable word problem to test whether the word equals the identity in the Artin monoid, and hence in the group.
@@
In addition to the submonoid membership problem, there are a range of other natural membership-type problems in groups that have been introduced and investigated including: rational subset membership, semigroup intersection, the group problem, and the identity problem (see below for definitions of these notions). The study of these decision problems has motivation coming from both mathematics and theoretical computer science, and is a burgeoning field of research with a lot of recent intensive activity; see e.g. the recent survey articles [Don23, Loh24] and also [Loh15]. For a helpful diagram showing how all these decision problems relate to each other we refer the reader to [Don23, Fig. 1]. In addition to the submonoid membership problem, we shall also consider these other decision problems listed above for Artin groups. We will show that our classification of Artin groups with decidable submonoid membership problem also gives a classification of Artin groups with decidable rational subset membership problem, and with decidable semigroup intersection problem; see Corollary 4.6.
@@
We do not currently know if this extends to also classify the Artin groups with decidable group problem or identity problem. In this direction in Theorem 4.5 we identify some families of Artin groups for which the identity problem and group problem are undecidable, and in the remaining cases of our classification we show that a related problem is undecidable: the fixed-target submonoid membership problem (that we shall define). These results are established by proving several new undecidability results for the right-angled Artin group $A(P_{4})$, where $P_{4}$ denotes the path graph on four vertices, in Section 3.
@@
In addition to presenting these algorithmic results on braid and Artin groups, another aim of this paper is to bring to the attention to those working on group and semigroup membership problems this method of embedding $A(P_{4})$ as an approach to proving undecidability results for groups that are known not to contain $F_{2}\times F_{2}$. Indeed, it was recently proved in [FGNB24, Corollary 6.4] that just embedding the trace monoid $T(P_{4})$ into a group is enough to yield undecidability of membership in rational subsets. We believe that there are likely many more natural situations when pairs of words cannot be embedded into a group, but $A(P_{4})$ or $T(P_{4})$ can, giving rise to undecidability results. This is, for example, exactly what happens in the case of one-relator groups and monoids; see [Gra20, FGNB24] and also the related work [MV24].
The main results of the article are the following:
  • We classify precisely when the submonoid (and rational subset) membership problem is decidable in braid groups (Theorem 3.1).
  • We show that the fixed-target submonoid membership problem is undecidable in $\mathbf{B}_{4}$ (Theorem 3.7).
  • We classify precisely when the submonoid membership problem is decidable in Artin groups in terms of the non-existence of certain forbidden induced subgraphs of the defining graph (Theorem 4.4).
  • We show that the same classification also applies for decidability of other decision problems (the rational subset membership and the semigroup intersection problem) in Artin groups (Corollary 4.6).
  • We identify some families of Artin groups for which the identity problem and group problem are undecidable (Theorem 4.5).
Acknowledgements
We wish to thank the anonymous referees for their careful reading of the article and useful comments.
@@ 2.1. Monoid presentations and rational subsets
A subset of $A^{\ast}$ is called a language. A finite automaton $\operatorname{\mathcal{A}}$ (over the alphabet $A$) consists of a finite set $Q$ of states, a finite set $\Delta\subseteq Q\times A\times Q$ of transitions, a distinguished start state $q_{0}\in Q$, and a set $F\subseteq Q$ of final states. We will consider $\operatorname{\mathcal{A}}$ as a directed graph, with vertex set $Q$ and an edge $q\xrightarrow{a}q^{\prime}$ whenever $(q,a,q^{\prime})\in\Delta$. The language accepted by $\operatorname{\mathcal{A}}$, denoted $L(\operatorname{\mathcal{A}})$, is the set of all words $w\equiv a_{1}a_{2}\cdots a_{n}$, where $x_{i}\in A$, for which there exist $q_{1},\dots,q_{n}\in Q$ such that for all $1\leq i\leq n$, we have $(q_{i-1}\xrightarrow{a_{i}}q_{i})\in\Delta$, i.e. such that $w$ labels a directed path from $q_{0}$ to some state in $F$. Any language $R\subseteq A^{\ast}$ such that $R=L(\operatorname{\mathcal{A}})$ for some finite automaton $\operatorname{\mathcal{A}}$ is called regular. To simplify some technical steps, we will also allow $\varepsilon$-transitions, being edges of the form $q_{i}\xrightarrow{\varepsilon}q_{j}$. This means that we can pass from state $q_{i}$ to $q_{j}$ without reading any letter. It is well-known that the class of regular languages is the same as the class of languages accepted by finite automata allowing $\varepsilon$-transitions.
@@ 2.2. Membership problems
There are many actively studied decision problems which fall under these umbrella terms. Let $G$ be a group generated by a finite set $A$, with a surjective homomorphism $\pi\colon A^{\ast}\to G$. The (uniform) submonoid membership problem for $G$ asks, on input a finite set $X\subseteq A^{\ast}$ of words, and a word $w\in A^{\ast}$, whether or not $\pi(w)\in\operatorname{Mon}(X)$. The non-uniform submonoid membership is the same problem, except the finite subset $X$ is considered fixed and not part of the input; similarly, the fixed-target submonoid membership problem fixes the word $w$, and only has $X$ as part of the input. This problem will thus be referred to as the “fixed-target submonoid membership problem for $w$ in MM”. All the above terminology is retained if instead of submonoids we consider membership in subsemigroups $\operatorname{Sgp}(X)$. A particular case of the fixed-target subsemigroup problem is when the fixed target is $1$, i.e. when we are given a set $X$ and asked to decide if $1\in\operatorname{Sgp}(X)$. This problem is known in the literature, e.g. [Don23], as the identity problem [^1] The reader is warned that the terminology “identity problem” is somewhat unfortunate, as this is also the classical name for the word problem, used e.g. by Dehn [Deh11].. Analogously, the (uniform) rational subset membership problem for $G$ asks, on input a finite automaton $\operatorname{\mathcal{A}}$ over $A$ and a word $w\in A^{\ast}$, whether or not $\pi(w)\in\pi(L(\operatorname{\mathcal{A}}))$. The non-uniform version of this problem is the same, except the automaton $\operatorname{\mathcal{A}}$ is fixed, and not part of the input. In principle, one could consider the fixed-target rational subset membership problem, fixing a target element $w$, and leaving $\operatorname{\mathcal{A}}$ as the only input. However, this can easily seen to be equivalent to the (uniform) rational subset membership problem; we thank C. Bodart for pointing out this fact to us.
@@
Another important membership-type problem that will be studied herein is the subsemigroup intersection problem, which asks, on input two finite subsets $X_{1},X_{2}\subseteq A^{\ast}$, whether or not $\operatorname{Sgp}(X_{1})\cap\operatorname{Sgp}(X_{2})=\varnothing$ in $G$. This also has more and less uniform, as well as fixed-target, variants. For example, one might fix one of the subsets $X_{i}$ (making the problem less uniform), or ask if some fixed element sits in the intersection (making the problem of a fixed-target type). We remark that if a group $G$ has decidable rational subset membership problem, then we can also solve the identity problem and the subsemigroup intersection problem in $G$. Indeed, the first is immediate, and the latter follows from the fact that $S_{1}S_{2}^{-1}$ is a rational subset of $G$ for any two finitely generated subsemigroups $S_{1},S_{2}$. But $S_{1}\cap S_{2}\neq\varnothing$ if and only if $1\in S_{1}S_{2}^{-1}$, and since we decide the latter, we can also decide the former. Finally, another decision problem is the group problem, which takes as input a finite set $X$ of words and decides whether or not $\operatorname{Sgp}(X)$ is a subgroup of $G$.
@@
The decision problems discussed in this section behave well when passing from a finitely generated group to a finitely generated subgroup. More precisely, let $G$ be a finitely generated group and $H\leq G$ a finitely generated subgroup, and suppose that $\mathcal{P}$ is any of the following decision problems: submonoid membership problem, subgroup membership problem, rational subset membership problem, fixed-target submonoid membership problem with a fixed target element $\gamma$, subsemigroup intersection problem, group problem, and identity problem. Then it is straightforward to see that if $\mathcal{P}$ is decidable in $G$ then $\mathcal{P}$ is also decidable in $H$. This is an important tool for showing various problems are undecidable in a group, by embedding a suitable group for which we know the problem is undecidable.
@@ 2.3. Artin and braid groups
$\underbrace{abab\ldots}_{\mbox{$m$-factors}}=\underbrace{baba\ldots}_{\mbox{$m$-factors}}$
@@
with the first set of relations being taken for all $1\leq i<n-1$, and the second set only being taken for those $i,j$ such that $|i-j|\geq 2$. There is a natural pictorial description of any element (braid) of a braid group, which also gives rise to a simple way to multiply braids together; see [KT08, §1.2.2]. Clearly the $n$-strand braid groups $\mathbf{B}_{n}$ are all Artin groups. For example, the braid group $\mathbf{B}_{4}$ is the Artin group $A(\Gamma)$ where $\Gamma$ is a triangle with edges labelled $(2,3,3)$.
@@ 3. Braid groups and right-angled Artin groups
Theorem 3.1 will be proved by applying several results from the literature on right-angled Artin groups and their embeddings into braid groups. First, we have the following result of Lohrey & Steinberg in which they classify the right-angled Artin groups with decidable submonoid, and rational subset, membership problem.
@@
A right-angled Artin group $A(\Gamma)$ has decidable submonoid membership problem if and only $\Gamma$ does not contain $C_{4}$ or $P_{4}$ as an induced subgraph. Furthermore, if $\Gamma$ does not contain $C_{4}$ or $P_{4}$, then $A(\Gamma)$ has decidable rational subset membership problem, while if $\Gamma$ does contain $C_{4}$ or $P_{4}$ then $A(\Gamma)$ contains a fixed finitely generated submonoid in which membership is undecidable.
Here, and throughout this present article, $C_{4}$ denotes the cycle graph on four vertices, and $P_{4}$ denotes the path graph on four vertices. As mentioned in the introduction, there is no embedding of pairs of words, i.e. the trace monoid $T(C_{4})$, into $\mathbf{B}_{4}$, and hence the group $A(C_{4})\cong F_{2}\times F_{2}$ also does not embed into the braid group $\mathbf{B}_{4}$. However, the following result of Droms, Lewin & Servatius shows that the group $A(P_{4})$ does embed into $\mathbf{B}_{4}$.
@@ 3.1. A connection between rational subsets and submonoids
which is a finite set of size $|Q|$. We fix a bijection between the two sets $Q$ and $\widetilde{Q}$ such that $q_{0}\mapsto x$ and $q_{f}\mapsto y$. For $q\in Q$, let $\widetilde{q}$ denote the image of $q$ in $\widetilde{Q}$. Clearly, $\widetilde{Q}$ generates a free subgroup of $F$ of rank $|Q|$.
@@
$\Delta=\{\widetilde{p}\sigma\widetilde{q}^{-1}\mid p\xrightarrow{\sigma}q\text{ is a transition in $\operatorname{\mathcal{A}}$}\}$
@@
Thus, assume $n>1$, and that the claim holds for n1n-1. There are several subcases to consider. First, suppose $q_{i}=p_{i+1}$ for some $i$. In this case, we can cancel these terms, and since we then have a path $p_{i}\xrightarrow{v_{i}v_{i+1}}q_{i+1}$, we can apply induction. On the other hand, if $q_{i}=p_{i}$ and $v_{i}=1$ for some $1\leq i\leq n$, then we can of course eliminate $\widetilde{p_{i}}v_{i}\widetilde{q_{i}}^{-1}$ from (5) and obtain a shorter expression, and we are done by induction. Thus, the last case to consider is that $q_{i}\neq p_{i+1}$ and that $p_{i}=q_{i}$ implies $v_{i}\neq 1$ for all $1\leq i\leq n$. Said otherwise, this says that the right-hand side of (5) is in normal form, and hence, since the left-hand side is also in normal form, we must have $n=M$, $v_{i}=1$ in $G$, as well as $\widetilde{p_{i}}=x$ and $\widetilde{q_{i}}=y$ for all $1\leq i\leq n$. In particular, $p_{1}=q_{0}$ and $q_{1}=q_{f}$, and since $p_{1}\xrightarrow{v_{1}}q_{1}$, we have $q_{0}\xrightarrow{1}q_{f}$. In other words, we have $1\in L(\operatorname{\mathcal{A}})$, as required. ∎
@@
[…]
@@
Let us now prove the equivalence (4), which will complete the proof of part (ii) of the lemma, and thereby also the entire lemma. If $\pi(xy^{-1})\in\operatorname{Mon}(\Delta)$, then clearly as $xy^{-1}\neq 1$ in $G\ast F$, we have that $\pi(xy^{-1})\in\operatorname{Sgp}(\Delta)$, and hence $\operatorname{Sgp}(xy^{-1})\cap\operatorname{Sgp}(\Delta)\neq\varnothing$ as required. Suppose next, for the converse, that $\operatorname{Sgp}(\Delta)\cap\operatorname{Sgp}(xy^{-1})\neq\varnothing$. Then some power of $xy^{-1}$ lies in $\operatorname{Sgp}(\Delta)$, i.e.
@@
in $G\ast F$, where the $p_{i}\xrightarrow{a_{i}}q_{i}$ are paths in $\operatorname{\mathcal{A}}$. By the above claim, we have $1\in\pi(L(\mathcal{A}))$. Hence, by (3), we have that $\pi(xy^{-1})\in\operatorname{Mon}(\Delta)$. This completes the proof of (4), and hence also the proof of the lemma. ∎